# references.bib

### v4

@@ 6.1 Fourier transform of nilpotent Lie group $C^{\ast}$-algebras
In this chapter, we focus on verifying the compactness condition outlined in Definition 4.5.1, i.e., let $e$ be an element in the multiplier algebra of the $C^{\ast}(Gr(H)_{x})$-module $E=C^{\ast}(Gr(H)_{x})$, $e\in M(C^{\ast}(Gr(H)_{x}))$, we want to show whether or not $e$ lies in $\mathcal{K}(E)=E$.
@@
Recall that for a differential operator $P$, let $F$ be the ordinary cosymbol of $P$ we discussed in Remark 1.4.6. We can take the Fourier transformation of $F\in M(C^{\ast}(TM))$ to the symbol $f=\widehat{F}$ in $M(C_{0}(T^{\ast}M))=C_{b}(T^{\ast}M)$. Let us describe an idea to show the compactness of $F$ roughly as follows: We show that, when $\xi$ converges to infinity, $f(x,\xi)$ converges to 00, and thus, it lies in $C_{0}(T^{\ast}M)$ and then the original cosymbol operator $F$ is a compact multiplier of $C^{\ast}(TM)$. See the commutative diagram below:
@@
In case of a differential operator $P$ on the vector bundle $E$, the ordinary symbol can similarly be understood as an element in $C_{b}(M;p^{\ast}E)\cong C_{b}(T^{\ast}M)\otimes_{C(M)}C(M;E)$, with a compact symbol in the subspace $C(M;p^{\ast}E)\cong C_{0}(T^{\ast}M)\otimes_{C(M)}C(M;E)$.
@@
For an element $e\in M(C^{\ast}(\mathcal{G}))$, we show its Fourier transform $\mathcal{F}(e)\in\mathcal{F}(M(C^{\ast}(\mathcal{G})))$ lies in $\mathcal{F}(C^{\ast}(\mathcal{G}))$ by using the $C^{\ast}$-structure of the resulting space of Fourier transform. Then, $e\in\mathcal{K}(C^{\ast}(\mathcal{G}))=C^{\ast}(\mathcal{G})$. A parallel diagram illustrates the relationship:
@@
We can apply $\mathcal{G}$ to be $Gr(H)_{x}$ for each $x$, which is nilpotent. In this case, we can verify the compactness of a multiplier on $C^{\ast}(\mathcal{G})$ for a nilpotent Lie group $\mathcal{G}$ by using the theory regarding Fourier transform of nilpotent Lie group $C^{\ast}$-algebra, which is treated in [beltictua2017fourier].
@@ 6.2 Application to the multiplier algebra
The multiplier algebra of $C_{0}$ functions on a topological space $X$ with values in a certain $C^{\ast}$-algebra can be calculated as follows.
@@
Let $A$ be a $C^{\ast}$-algebra. Let $M(A)$ be its multiplier $C^{\ast}$-algebra. The strict topology on $M(A)$ is given by the convergence condition:
@@
${M}(C_{0}(X,A))=C_{b}(X,M(A)_{\beta}),$
where $M(A)_{\beta}$ is the multiplier algebra of $A$ with the strict topology.
@@
A multiplier of $M(B^{\ast}(S))$ is an element of the operator field $\psi$ over $S$ that satisfies:
@@
Take a multiplier $F\in M(C^{\ast}(\mathcal{G}))$, $F:C^{\ast}(\mathcal{G})\to C^{\ast}(\mathcal{G})$. For each representation $\pi\in\widehat{\mathcal{G}}$, $\pi$ extends to a representation of $C^{\ast}(\mathcal{G})$; it induces a map $\pi(F):=F\otimes_{\pi}1$ from $C^{\ast}(\mathcal{G})\otimes_{\pi}\mathcal{H}_{\pi}\cong\mathcal{H_{\pi}}$ to itself. Then by Definition 6.1.1, we find that $\{\pi(F)\}_{\pi\in\widehat{\mathcal{G}}}$ is the Fourier transform of an element $F$ in the multiplier algebra $M(C^{\ast}(\mathcal{G}))$.
@@
A multiplier $Q\in M(E_{x}\otimes C^{\ast}(Gr(H)_{x}))$ is compact, i.e., $Q\in\text{End}(E_{x})\otimes C^{\ast}(Gr(H))_{x}$ if and only if the operator field
@@
    In general cases, the direct decomposition of representations (aligned with coadjoint orbits) into transverse and leaf-wise components, similar to the partitioning of cotangent vectors, presents significant challenges. This complexity leads to the difficulties involved in addressing transversal H-ellipticity. Nevertheless, the application of Fourier transform methodologies facilitates the transformation of operators to actions on Hilbert spaces, thereby simplifying the task to prove the compactness of these operators within each Hilbert space. This approach allows for the partitioning of representing Hilbert spaces into distinct segments, facilitating a separate examination of the transverse and leaf-wise directional elements.
@@ A.1 Other ideas for the definition of transversal H-ellipticity
Alternative definitions of transversal H-ellipticity can be considered. All the definitions are motivated by some characteristics of H-ellipticity or transversal ellipticity. However, these definitions are not sufficient to calculate in $KK$-theory and it’s hard to show the equivalent or inclusion relations between any two of these different conditions, as the $C^{\ast}$-algebra $C^{\ast}(Gr(H))$ is non-commutative and we cannot regularly decompose $gr(H)_{x}$ elements into transversal and leaf-wise directions. These definitions only provide other possible directions to analyze.
@@ A.1.1 The main definition
Let $M$ be a compact filtered Riemannian manifold and $G$ be a Lie group acting on $M$ isometrically and preserving the filtered structure. We call a pseudodifferential operator of 00-order $P$ is transversally H-elliptic if at each point $x\in M$, the multiplier term
$(1-\sigma_{H}^{2}(P,x))(1+\sigma_{H}(\Delta_{G},x))^{-1}\in M(E\otimes_{C(M)}C^{\ast}(Gr(H)_{x}))$
@@ A.1.2 Definition from Atiyah’s approach
The second definition is stated as follows:
@@
Inspired by Atiyah’s approach [atiyah2006elliptic], the second definition (Definition B) focuses on the injectivity of symbols for the operator $P$ and its adjoint $P^{\ast}$ when combined with a leaf-wise elliptic Laplacian operator. This perspective allows for the analytical index to be defined as a distribution over the group $G$, employing Atiyah’s method.
@@ A.1.3 An intuitive definition
In Chapter 4, we have defined the bundle isomorphism $\psi:TM\to gr(H)$, which maps a vector $v\in T_{x}M$ decomposed as $v=\sum_{i}v^{i}$ to $\psi(v)=\sum_{i}[v^{i}]\in gr(H)_{x}$, with each $v^{i}\in H^{i}_{x}$ and all $v^{i}$’s are orthogonal to each other. We obtain the adjoint map $\psi^{\ast}:gr(H)^{\ast}\to T^{\ast}M$. The image of transversal part $T^{\ast}_{G,x}M$ by $(\psi^{\ast})^{-1}$ is denoted as $U^{\ast}_{G,x}\subseteq gr(H)_{x}^{\ast}$. We define
@@
The third definition is stated as follows:
@@
$P$ is called transversally H-elliptic if for all $x\in M$ $\sigma(P,x,\xi)$ is bijective for $\pi\in O^{\ast}_{G,x}$, i.e., $\sigma(P,x,\pi_{\xi})$ is bijective if the coadjoint orbit corresponding to $\pi_{\xi}$ contains at least one point of $T^{\ast}_{G}M$.
The third definition (Definition C) is the most straightforward one, directly aligning with the foundational concept of transversal ellipticity. However, its practical application within $KK$-theory is restricted by the lack of information on the representations out of $O^{\ast}_{G,x}$. There is also a problem that we cannot decompose coadjoint orbits of the nilpotent Lie group $Gr(H)_{x}$, similar to the decomposition of $T_{x}M$ to tangent part and transversal part, because $C^{\ast}(Gr(H))$ is non-commutative,
Adopting this third definition (Definition C) facilitates a construction parallel to that for transversally elliptic cases by defining a specialized $KK$-class
$[\mathcal{E}^{\prime}]\in KK_{G}(C^{\ast}(O_{G}),C_{0}(T^{\ast}_{G}M)).$
Define
$C^{\ast}(O_{G,x})=\frac{C^{\ast}(Gr(H)_{x})}{\bigcap_{\xi\in O_{G,x}}\ker\pi_{\xi}}.$
The $KK$-class $[\mathcal{E}]$ can be restricted to define a $KK$-class $[\mathcal{E}^{\prime}]\in KK_{G}(C^{\ast}(O_{G}),C_{0}(T^{\ast}_{G}M))$.
Thus, if we consider such transversally H-elliptic operators in Definition C with “good” symbol (which is defined in [berline1996chern]), we can typically define the symbol class $[\sigma^{\prime}_{H}(P)]\in KK_{G}(\mathbb{C},$ $C^{\ast}(O_{G}))$, and index class $[P]\in KK(C^{\ast}(G),\mathbb{C})$. Then $[\mathcal{E}^{\prime}]$ in a certain sense yields a restricted Connes-Thom isomorphism. If we define the Dolbeault operator $[\mathcal{D}_{M}^{\prime}]\in K^{0}(C^{\ast}(G,C_{0}(T^{\ast}_{G}M)))$ in transversally direction similarly, we obtain a formula that
$[P]=j^{G}([\sigma^{\prime}_{H}(P)]\otimes_{C^{\ast}(O_{G})}[\mathcal{E}^{\prime}])\otimes_{C^{\ast}(G,C_{0}(T^{\ast}_{G}M))}[\mathcal{D}_{M}^{\prime}].$

### v5

@@ 6.1 Fourier transform of nilpotent Lie group $C^{\ast}$-algebras
Recall that for a differential operator $P$, let $F$ be the ordinary cosymbol of $P$ we discussed in Remark 1.4.7. We can take the Fourier transformation of $F\in\mathcal{M}(C^{\ast}(TM))$ to the symbol $f=\widehat{F}$ in $\mathcal{M}(C_{0}(T^{\ast}M))=C_{b}(T^{\ast}M)$. Let us describe an idea to show the compactness of $F$ roughly as follows: We show that, when $\xi$ converges to infinity, $f(x,\xi)$ converges to 00, and thus, it lies in $C_{0}(T^{\ast}M)$ and then the original cosymbol operator $F$ is a compact multiplier of $C^{\ast}(TM)$. See the commutative diagram below:
@@
In case of a differential operator $P$ on the vector bundle $E$, the ordinary symbol can similarly be understood as an element in $C_{b}(M;p^{\ast}E)\cong C_{b}(T^{\ast}M)\otimes_{C_{0}(M)}C_{0}(M;E)$, with a compact symbol in the subspace $C_{0}(M;p^{\ast}E)\cong C_{0}(T^{\ast}M)\otimes_{C_{0}(M)}C_{0}(M;E)$.
@@
For an element $e\in\mathcal{M}(C^{\ast}(\mathcal{G}))$, we show its Fourier transform $\mathcal{F}(e)\in\mathcal{F}(\mathcal{M}(C^{\ast}(\mathcal{G})))$ lies in $\mathcal{F}(C^{\ast}(\mathcal{G}))$ by using the $C^{\ast}$-structure of the resulting space of Fourier transform. Then, $e\in\mathcal{K}(C^{\ast}(\mathcal{G}))=C^{\ast}(\mathcal{G})$. A parallel diagram illustrates the relationship:
@@
We can assign $\mathcal{G}$ to be $Gr(H)_{x}$ for each $x$, which is nilpotent. In this case, we can verify the compactness of a multiplier on $C^{\ast}(\mathcal{G})$ for a nilpotent Lie group $\mathcal{G}$ by using the theory regarding Fourier transform of nilpotent Lie group $C^{\ast}$-algebra, which has been treated in [beltictua2017fourier].
@@ 6.2 Application to the multiplier algebra
By the main idea at the beginning of Section 6.1 and Corollary 6.1.7, it suffices to analyze the multiplier algebra of $B^{\ast}(\widehat{\mathcal{G}})$. Then we can give an answer for the Question 6.0.1.
The multiplier algebra of $C_{0}$ functions on a topological space $M$ with values in a certain $C^{\ast}$-algebra can be calculated as follows.
@@
Let $A$ be a $C^{\ast}$-algebra. Let $\mathcal{M}(A)$ be its multiplier $C^{\ast}$-algebra. The strict topology on $\mathcal{M}(A)$ is given by the convergence condition:
@@
${M}(C_{0}(M,A))=C_{b}(M,\mathcal{M}(A)_{\beta}),$
where $\mathcal{M}(A)_{\beta}$ is the multiplier algebra of $A$ with the strict topology.
@@
A multiplier of $\mathcal{M}(B^{\ast}(S))$ is an element of the operator field $\psi$ over $S$ that satisfies:
@@
Take a multiplier $F\in\mathcal{M}(C^{\ast}(\mathcal{G}))$, $F:C^{\ast}(\mathcal{G})\to C^{\ast}(\mathcal{G})$. For each representation $\pi\in\widehat{\mathcal{G}}$, $\pi$ extends to a representation of $C^{\ast}(\mathcal{G})$; it induces a map $\pi(F):=F\otimes_{\pi}1$ from $C^{\ast}(\mathcal{G})\otimes_{\pi}\mathcal{H}_{\pi}\cong\mathcal{H_{\pi}}$ to itself. Then by Definition 6.1.1, we find that $\{\pi(F)\}_{\pi\in\widehat{\mathcal{G}}}$ is the Fourier transform of an element $F$ in the multiplier algebra $\mathcal{M}(C^{\ast}(\mathcal{G}))$.
@@
A multiplier $Q\in\mathcal{M}(E_{x}\otimes C^{\ast}(Gr(H)_{x}))$ is compact, i.e., $Q\in\text{End}(E_{x})\otimes C^{\ast}(Gr(H))_{x}$ if and only if the operator field
@@
    In general cases, the direct decomposition of representations (aligned with coadjoint orbits) into transverse and leaf-wise components, similar to the partitioning of cotangent vectors, presents significant challenges. This complexity leads to difficulties in addressing transversal H-ellipticity. Nevertheless, the application of Fourier transform methodologies facilitates the transformation of operators to actions on Hilbert spaces, thereby simplifying the task to prove the compactness of these operators within each Hilbert space. This approach allows for the partitioning of representing Hilbert spaces into distinct segments, facilitating a separate examination of the transverse and leaf-wise directional elements.
@@ A.1 Other ideas for the definition of transversal H-ellipticity
Alternative definitions of transversal H-ellipticity can also be considered. All the definitions are motivated by some characteristics of H-ellipticity or transversal ellipticity. However, these definitions are not sufficient to calculate in $KK$-theory and it’s hard to show the equivalent or inclusion relations between any two of these different conditions, because the $C^{\ast}$-algebra $C^{\ast}(Gr(H))$ is non-commutative and we cannot regularly decompose $gr(H)_{x}$ elements into transversal and leaf-wise directions. These definitions only provide other possible directions to analyze.
@@ A.1.1 The main definition
Let $M$ be a complete filtered Riemannian manifold and $G$ be a Lie group acting on $M$ isometrically and preserving the filtered structure. We call a pseudodifferential operator of 00-order $P$ is transversally H-elliptic if at each point $x\in M$, the multiplier term
$(1-\sigma_{H}^{2}(P,x))(1+\sigma_{H}(\Delta_{G},x))^{-1}\in\mathcal{M}(E_{x}\otimes C^{\ast}(Gr(H)_{x}))$
@@ A.1.2 Definition from Atiyah’s approach
The second possible definition is stated as follows:
@@
Inspired by Atiyah’s approach [atiyah2006elliptic], the second definition (Definition B) focuses on the injectivity of symbols for the operator $P$ and its adjoint $P^{\ast}$ when combined with the leaf-wise elliptic Laplacian operator. This perspective allows for the analytical index to be defined as a distribution over the group $G$, employing Atiyah’s method.
@@ A.1.3 An intuitive definition
In Chapter 4, we have defined the bundle isomorphism $\psi:TM\to gr(H)$, which maps a vector $v\in T_{x}M$ decomposed as $v=\sum_{i}v^{i}$ to $\psi(v)=\sum_{i}[v^{i}]\in gr(H)_{x}$, with each $v^{i}\in H^{i}_{x}$ and all $v^{i}$’s are orthogonal to each other. We obtain the adjoint map $\psi^{\ast}:gr(H)^{\ast}\to T^{\ast}M$. The image of transversal part $T^{\ast}_{G,x}M$ by $(\psi^{\ast})^{-1}$ is denoted as $U^{\ast}_{G,x}\subseteq gr(H)_{x}^{\ast}$. Then, we define
@@
The third possible definition is stated as follows:
@@
$P$ is called transversally H-elliptic if for all $x\in M$, $\sigma(P,x,\xi)$ is bijective for $\pi\in O^{\ast}_{G,x}$, i.e., $\sigma(P,x,\pi_{\xi})$ is bijective if the coadjoint orbit corresponding to $\pi_{\xi}$ contains at least one point of $T^{\ast}_{G}M$.
The third definition (Definition C) is the most straightforward one, directly aligning with the foundational concept of transversal ellipticity. However, its practical application within $KK$-theory is restricted by the lack of information on the representations out of $O^{\ast}_{G,x}$. There is also a problem that we cannot decompose coadjoint orbits of the nilpotent Lie group $Gr(H)_{x}$, similar to the decomposition of $T_{x}M$ to tangent part and transversal part, because $C^{\ast}(Gr(H))$ is non-commutative.